# Diffusion differentiable resampling

### v1

@@ 4.4 Vision-based pendulum dynamics tracking
where $Z_{j}\coloneqq\begin{bmatrix}Z^{(1)}_{j}&Z^{(2)}_{j}\end{bmatrix}$ represents the state defined by the angle and angular velocity, $\Delta_{j}$ is the discretisation step, and $\zeta_{j}\sim\mathcal{N}(0,\sigma_{\zeta}^{2}\,\Delta_{j}\,I_{2})$ is the process noise. The true observation function $r$ generates $32\times 32$ greyscale images representing the pendulum state. We generate a single trajectory of observations $Y_{0:J}$ over steps $j=0,1,\ldots,256$. Crucially, in contrast to the previous example in Section 4.3, we here assume that both the underlying dynamics and the observation model are unknown, and we parametrise both the state transition $Z_{j}=f_{\theta}(Z_{j-1},\zeta_{j})$, and the decoding $r\approx r_{\phi}$, by neural networks. Parameters $\theta$ and $\phi$ are learnt by minimising the negative log-likelihood estimated by the particle filter with resampling. We evaluate the learnt models by comparing the average SSIM and PSNR of their predicted image sequences. All results are averaged over 5 independent runs, using $N=32$ particles.
Our results show that diffusion resampling enables learning meaningful latent representations and making high-fidelity predictions, achieving a performance that is at least comparable to state-of-the-art differentiable resampling baselines. The comparative results in Figure 5 show that, among the best-performing configurations of each resampling class, diffusion resampling yields the highest mean SSIM and PSNR. Figure 6 shows the mean image-sequence reconstruction from a pendulum dynamics model and decoder learnt with diffusion resampling, illustrating that diffusion resampling facilitates learning of complex SSMs in high-dimensional and complex settings. For further details, see Appendix J.
Figure 5: Mean prediction SSIM and PSNR for the best (by mean) model configuration of each resampler. Individual runs are shown as scatter points.
Figure 6: Qualitative comparison of the learnt pendulum dynamics (trained with diffusion resampling until convergence) against the ground truth. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom).
@@ 5 Related work
A prior work closely related to ours is that of Wan2025diffpf, who also leverages a reverse diffusion process for differentiable resampling within the SMC framework. While empirically powerful, their approach relies on a learnt conditional diffusion model, which introduces irreducible bias, breaks consistency guarantees, and adds substantial computational cost. In addition, their method introduces a further challenge in that the resampling gradient should also be propagated through the training process. In contrast, our method focuses on a statistically grounded diffusion resampling scheme without training, and is more informative.
There are also connections to other, less central, lines of work. For instance, the ensemble score estimator is related to the kernel projection used in Stein variational gradient flow (Liu2017stein). The controlled gain term in Yang2013 also plays a role analogous to resampling in a continuous-time limit, but involves a computationally demanding PDE-solving step. While there are interesting connections to our work, the approach focuses on continuous-time flow filtering (cf. Kang2025) and lead to algorithms that are no longer standard SMC.
@@ 6 Conclusion
In this paper we have proposed a new differentiable-by-construction, computationally efficient, and informative resampling method built on diffusion models. We have established the consistency result in the number of samples and explicitly quantified the error bound. Our experiments empirically verify that the proposed method significantly outperforms the state-of-the-art differentiable resampling methods for resampling, filtering, and parameter estimation of state-space models.
Limitations and future work. We observe that propagating gradients through the diffusion resampling can be numerically sensitive to the SDE solver. This could be potentially mitigated by, e.g., reversible adjoint Heun’s method (Kidger2021gradient) or related variants. We have also mentioned the possibility to speed up the computation via a Gibbs sampler which warrants future development. It is also important to further tighten the theoretical error bound by analysing the conditions when the resampling error becomes optimal. We finally note that the ensemble score may not be the only choice for constructing the diffusion resampling, as long as the strategy in Remark 1 retains.
@@ 3 $\alpha_{i}(x,t)\leftarrow\alpha_{i}(x,t)\,/\,\sum_{j=1}^{N}\alpha_{j}(x,t)$
$\sum_{i=1}^{N}\alpha_{i}(x,t)\,\biggl(-\frac{x-m_{t}(X_{i})}{V_{t}}\biggr)$
@@ Appendix B Proof of Proposition 1
where $h^{\prime}_{\tau}$ stands for the derivative of $h_{\tau}$. We obtain
$\begin{split}\lvert h_{t}\rvert^{2}&=\lvert h_{0}\rvert^{2}+2\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(\widetilde{U}(\tau))-\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(U(\tau))+2\,\operatorname{\nabla\!}\log p_{T-\tau}(U(\tau))-s_{N}(\widetilde{U}(\tau),T-\tau)\bigr\rangle\mathop{}\!\mathrm{d}\tau\\ &=\lvert h_{0}\rvert^{2}+2\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(\widetilde{U}(\tau))-\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(U(\tau))\bigr\rangle\mathop{}\!\mathrm{d}\tau\\ &\qquad\qquad+4\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log p_{T-\tau}(U(\tau))-\operatorname{\nabla\!}\log p_{T-\tau}(\widetilde{U}(\tau))\bigr\rangle+\bigl\langle h_{\tau},\operatorname{\nabla\!}\log p_{T-\tau}(\widetilde{U}(\tau))-s_{N}(\widetilde{U}(\tau),T-\tau)\bigr\rangle\mathop{}\!\mathrm{d}\tau\end{split}$
@@
$\begin{split}\lvert h_{t}\rvert^{2}&\leq\lvert h_{0}\rvert^{2}+2\,b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\int^{t}_{0}\bigl\lvert h_{\tau}\bigr\rvert^{2}\mathop{}\!\mathrm{d}\tau+4\,b^{2}\int^{t}_{0}\lvert h_{t}\rvert\,\frac{C_{e}(T-\tau)}{N^{r}}\mathop{}\!\mathrm{d}\tau\end{split}$
Applying (non-linear) Grönwall inequality (Dragomir2003, Theorem 21) we finally obtain
$\lvert h_{t}\rvert\leq\lvert h_{0}\rvert\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}+2\,b^{2}\,N^{-r}\int^{t}_{0}C_{e}(T-\tau)\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,(t-\tau)}\mathop{}\!\mathrm{d}\tau.$ (19)
This directly implies an upper bound for the 1-Wasserstein distance between the measures of $U(t)\sim q_{t}$ and $\widetilde{U}(t)\sim\widetilde{q}_{t}$:
$\mathsf{W}_{1}(\widetilde{q}_{t},q_{t})\leq\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})+2\,b^{2}\,N^{-r}\overline{C}_{e}(t,T),$ (20)
@@ Appendix C Proof of Corollary 1
The error $\mathsf{W}_{1}(p_{t},\pi_{\mathrm{ref}})\to 0$ geometrically fast as $t\to\infty$, that is,
$\mathsf{W}_{1}(p_{T},\pi_{\mathrm{ref}})\leq\mathrm{e}^{-b^{2}\,C_{\mathrm{ref}}\,T}\,\mathsf{W}_{1}(\pi,\pi_{\mathrm{ref}}),$ (22)
where $C_{\mathrm{ref}}$ is the convex constant in Corollary 1. See, for instance, Dalalyan2017 for the proof. Now choose $N^{r-c}=\overline{C}_{e}(t,T)$ for constant $0<c<r$, and substitute Equation (22) into Equation (20) we get
$\mathsf{W}_{1}(\widetilde{q}_{t},q_{t})\leq 2\,b^{2}\,N^{-c}+\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t-b^{2}\,C_{\mathrm{ref}}\,T}\,\mathsf{W}_{1}(\pi,\pi_{\mathrm{ref}}).$ (23)
@@
Due to the positivity of $C_{e}$, a sufficient condition for the derivative above being positive is $T^{\prime}(t)>0$ and $T^{\prime}(t)\leq 1$. The set of such sequence is not empty, and a trivial example is $T(t)=t+\delta$ for some parameter $\delta>0$, which is clearly independent of any of the system components. Moreover, the exponent $(C_{\mathrm{ref}}-2\,C_{p})\,t-C_{\mathrm{ref}}\,T(t)$ is always negative.
Note that the factor $N^{-c}$ is likely suboptimal since it can not exceed the importance sampling factor $r$. This is intuitive, because we are essentially asking $N$ to compensate the SDE accumulation error. It is possible to relax from this and choose $c\geq r$. This in turn means that we need assumptions on $C_{e}$ to ensure that $\overline{C}_{e}(t,T(t))$ is non-increasing in $t$, in contrast to what we have aimed for being increasing which is mild. However, such settings will make Assumption 2 more difficult to verify in practice. Since making mild and practically checkable assumptions on score approximation remains an open discussion (see, e.g., DeBortoli2022convergence; Chen2023sampling), we opt for a suboptimal bound under assumptions that are simple and easy-to-verify (n.b., we have only assumed $C_{e}$ to be a positive non-increasing function without specifying any rate).
The result essentially reveals how we can leverage information from $\pi_{\mathrm{ref}}$ to achieve for lower resampling error. Commonly used resampling schemes assume that we only have access to the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$. On the contrary, diffusion resampling additionally assumes a sampler for $\pi_{\mathrm{ref}}$ that approximates $\pi$ yielding more information. The diffusion resampling effectively fuses the samples from $\{(w_{i},X_{i})\}_{i=1}^{N}$ and $\pi_{\mathrm{ref}}$ to obtain the re-samples, with the constant $b$ indicating how much we use $\pi_{\mathrm{ref}}$. With $b=0$, it means that we completely trust the samples from $\pi_{\mathrm{ref}}$, and this is indeed optimal when $\pi_{\mathrm{ref}}=\pi$.
@@ Appendix E Common experiment settings
Implementations are based on JAX (Jax2018github). When applying a Wasserstein distance to quantify the quality of approximate samples, we use the earth-moving cost function denoted by $\mathsf{W}_{1}$, approximated by a sliced Wasserstein distance with 1,000 projections. We directly use the implementation by OTT-JAX (Cuturi2022OTT). For neural network implementations, we use Flax (Flax2020github).
@@
  • Four SDE integrators. The commonly used Euler–Maruyama, and the two exponential integrators by Jentzen2009 and Lord2004, see Section 2.3 for formulae, and Tweedie’s formula (Ho2020, essentially DDPM). They are called by EM, Jentzen–Kloeden, Lord–Rougemont, and Tweedie, respectively in the later context.
  • SDE and probability flow. The resampling SDE in Equation 4 can be solved also with a probability flow ODE (Song2021scorebased). We test both SDE and ODE versions.
@@
The seminal work by Jang2017categorical did not formulate how the Gumbel trick can be applied for resampling. We here make it explicit, and the gist is to approximate the discrete indexing of multinomial resampling by a matrix-vector multiplication which is differentiable. Recall our weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}$, and define a matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ aggregating all the $N$ samples. For each $i$, draw $N$ independent samples $\{u_{i,j}\sim\mathrm{Uniform}[0,1]\}_{j=1}^{N}$ and compute $g_{i,j}=-\log\log u_{i,j}$. Then, the $i$-th re-sample is
@@
where $S_{i}\in\mathbb{R}^{1\times N}$ is a Softmax vector with elements $S_{i,j}=\overline{S}_{i,j}\,/\,\sum_{j=1}^{N}\overline{S}_{i,j}$, where $\overline{S}_{i,j}=\exp((\log w_{j}+g_{i,j})\,/\,\tau)$. This is independently repeated for $i=1,2,\ldots,N$ to generate the re-samples $\{X^{*}_{i}\}_{i=1}^{N}$. Upon the limit $\tau\to 0$, this recovers multinomial resampling, but the variance of the gradient estimation will grow too.
@@
The approach by Karkus2018 view the samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ as a discrete distribution with probability given by their normalised weights. Denote this discrete distribution by $\pi^{\mathrm{D}}$. Directly sampling from $\pi^{\mathrm{D}}$ using weighted random choice will result in undefined gradients. Soft resampling mitigates this issue by introducing another proposal (discrete) distribution $q^{\mathrm{D}}$ defined by $q^{\mathrm{D}}(X_{i})\coloneqq\alpha\,w_{i}+(1-\alpha)\,\frac{1}{N}$ for $i=1,2,\ldots,N$. Then, sampling from $\pi^{\mathrm{D}}$ can be achieved by an importance sampling as follows.

### v2

@@ 4.4 Vision-based pendulum dynamics tracking
where $Z_{j}\coloneqq\begin{bmatrix}Z^{(1)}_{j}&Z^{(2)}_{j}\end{bmatrix}$ represents the state defined by the angle and angular velocity, $\Delta_{j}$ is the discretisation step, and $\zeta_{j}\sim\mathcal{N}(0,\Delta_{j}\,\Lambda_{\zeta})$ is the process noise. The true observation function $r$ generates $32\times 32$ greyscale images representing the pendulum state. We generate a trajectory of observations $Y_{0:J}$ over steps $j=0,1,\ldots,256$. Crucially, in contrast to the previous example in Section 4.3, we here assume that both the underlying dynamics and the observation model are unknown, and we parametrise both the state transition, $Z_{j}=f_{\theta}(Z_{j-1},\zeta_{j})$, and the decoder, $r\approx r_{\phi}$, by neural networks. Parameters $\theta$ and $\phi$ are learnt by minimising the negative log-likelihood estimated by the particle filter with resampling. For completeness, we run experiments in two regimes: the first uses adaptive resampling under a tempered observation likelihood, and the second is a more challenging setting with resampling at nearly every filtering step. The learnt models are evaluated using the average structural similarity index (SSIM) and peak signal-to-noise ratio (PSNR) on their predicted image sequences.
Our results demonstrate that diffusion resampling integrates effectively into this complex, high-dimensional SMC learning pipeline, enabling stable optimisation and achieving competitive performance relative to state-of-the-art differentiable resampling baselines. Figure 6 shows an accurate and high-fidelity mean image sequence reconstructed from a latent pendulum dynamics model and decoder learnt with diffusion resampling embedded in the optimisation process. The comparative results in Figure 5 show that, among the strongest configurations of each resampling class, diffusion resampling achieves mean SSIM and PSNR at least comparable to the baselines. See Appendix J for further details and results.
Figure 5: Mean prediction SSIM and PSNR for the best (by mean) model configuration of each resampler in the first experiment setting. Individual runs are shown as scatter points.
Figure 6: Qualitative comparison of the learnt pendulum dynamics (trained end-to-end in the first experiment setting) against the ground truth. The ground truth (green) is overlaid with model predictions (purple). White pixels indicate perfect alignment, while coloured regions highlight positional discrepancies (e.g., phase lag). Snapshots are shown at eight evenly spaced time points over the 4 second simulation (read from left to right and top to bottom).
@@ 5 Related work
In concurrent work, Gourevitch et al. (2026) propose a differentiable reparameterisation of categorical distributions, constructed using stochastic interpolants. While our diffusion resampler similarly constitutes a differentiable relaxation of categorical sampling, the key departure is that the closed-form denoiser in their work is derived under one-hot encoded samples $\{X_{i}\}_{i=1}^{N}$, whereas we focus on samples in $\mathbb{R}^{d}$. The two approaches coincide when the samples are discrete variables and when formulated within the same diffusion model framework.
Another work closely related to ours is that by Wan and Zhao (2025), who also leverages a diffusion model for differentiable resampling within the SMC framework. While empirically powerful, their approach relies on a trained conditional diffusion model, which introduces bias, breaks consistency guarantees, and adds substantial computational cost. In addition, their method introduces a further challenge in that the resampling gradient should also be propagated through the diffusion training. In contrast, our method focuses on a statistically grounded diffusion resampling scheme without training, and is more informative.
@@ 6 Conclusion
In this paper we have proposed a new differentiable-by-construction, computationally efficient, and informative resampling method built on diffusion models. We have explicitly quantified a convergent error bound in the sample size and diffusion parameters. Our experiments verify that the proposed method outperforms the state-of-the-art differentiable resampling methods for filtering and parameter estimation of state-space models.
Limitations and future work. We observe that propagating gradients through the diffusion resampling can be numerically sensitive to the SDE solver. This could potentially be mitigated by, e.g., reversible adjoint Heun’s method (Kidger et al., 2021) or related variants. We have also mentioned the possibility to speed up the computation via a Gibbs sampler which warrants future development. We finally note that the ensemble score may not be the only choice for diffusion resampling, as long as the strategy in Remark 1 retains.
@@ 3 $\alpha_{i}(x,t)\leftarrow\alpha_{i}(x,t)\,/\,\sum_{j=1}^{N}\alpha_{j}(x,t)$
$\sum_{i=1}^{N}\alpha_{i}(x,t)\,V_{t}^{-1}\,\bigl(x-m_{t}(X_{i})\bigr)$
@@ Appendix B Proof of Proposition 1
where $h^{\prime}_{\tau}$ stands for the drift of $h_{\tau}$. We obtain
$\begin{split}\lvert h_{t}\rvert^{2}&=\lvert h_{0}\rvert^{2}+2\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(\widetilde{U}(\tau))-\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(U(\tau))+2\,\operatorname{\nabla\!}\log p_{T-\tau}(U(\tau))-2\,s_{N}(\widetilde{U}(\tau),T-\tau)\bigr\rangle\mathop{}\!\mathrm{d}\tau\\ &=\lvert h_{0}\rvert^{2}+2\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(\widetilde{U}(\tau))-\operatorname{\nabla\!}\log\pi_{\mathrm{ref}}(U(\tau))\bigr\rangle\mathop{}\!\mathrm{d}\tau\\ &\qquad\qquad+4\,b^{2}\int^{t}_{0}\bigl\langle h_{\tau},\operatorname{\nabla\!}\log p_{T-\tau}(U(\tau))-\operatorname{\nabla\!}\log p_{T-\tau}(\widetilde{U}(\tau))\bigr\rangle+\bigl\langle h_{\tau},\operatorname{\nabla\!}\log p_{T-\tau}(\widetilde{U}(\tau))-s_{N}(\widetilde{U}(\tau),T-\tau)\bigr\rangle\mathop{}\!\mathrm{d}\tau.\end{split}$
@@
$\begin{split}\operatorname{\mathbb{E}}\bigl[\lvert h_{t}\rvert^{2}\bigr]&\leq\operatorname{\mathbb{E}}\bigl[\lvert h_{0}\rvert^{2}\bigr]+2\,b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\int^{t}_{0}\operatorname{\mathbb{E}}\bigl[\lvert h_{\tau}\rvert^{2}\bigr]\mathop{}\!\mathrm{d}\tau+4\,b^{2}\int^{t}_{0}\operatorname{\mathbb{E}}\bigl[\lvert h_{\tau}\rvert^{2}\bigr]^{\frac{1}{2}}\,\frac{C_{e}(T-\tau)}{N^{r}}\mathop{}\!\mathrm{d}\tau,\end{split}$
where recall that the weighted samples and the SDE system are mutually independent. Applying (non-linear) Grönwall inequality (Dragomir, 2003, Theorem 21) we finally obtain
$\operatorname{\mathbb{E}}\bigl[\lvert h_{t}\rvert^{2}\bigr]\leq\operatorname{\mathbb{E}}\bigl[\lvert h_{0}\rvert^{2}\bigr]\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}+2\,b^{2}\,N^{-r}\int^{t}_{0}C_{e}(T-\tau)\,\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,(t-\tau)}\mathop{}\!\mathrm{d}\tau.$ (19)
This directly implies an upper bound for the 2-Wasserstein distance between the measures of $U(t)\sim q_{t}$ and $\widetilde{U}(t)\sim\widetilde{q}_{t}$:
$\mathsf{W}_{2}^{2}(\widetilde{q}_{t},q_{t})\leq\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t}\mathsf{W}_{2}^{2}(p_{T},\pi_{\mathrm{ref}})+2\,b^{2}\,N^{-r}\overline{C}_{e}(t,T),$ (20)
@@ Appendix C Proof of Corollary 1
The error $\mathsf{W}_{1}(p_{t},\pi_{\mathrm{ref}})\to 0$ geometrically fast as $t\to\infty$, more specifically,
$\mathsf{W}_{2}^{2}(p_{T},\pi_{\mathrm{ref}})\leq\mathrm{e}^{-2\,b^{2}\,C_{\mathrm{ref}}^{-}\,T}\,\mathsf{W}_{2}^{2}(\pi,\pi_{\mathrm{ref}}),$ (22)
where $C_{\mathrm{ref}}^{-}$ is the concave constant in Assumption 1. This is a classical result, see, for instance, Ambrosio et al. (2008) and von Renesse and Sturm (2005). Now choose $N^{r-c}=\overline{C}_{e}(t,T)$ for any constant $c$ such that $0<c<r$, and substitute Equation (22) into Equation (20) we get
$\mathsf{W}_{2}^{2}(\widetilde{q}_{t},q_{t})\leq 2\,b^{2}\,N^{-c}+\mathrm{e}^{b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t-2\,b^{2}\,C_{\mathrm{ref}}^{-}\,T}\,\mathsf{W}_{2}^{2}(\pi,\pi_{\mathrm{ref}}).$ (23)
@@
Due to the positivity of $C_{e}$, a sufficient condition for the derivative above being positive is $T^{\prime}(t)>0$ and $T^{\prime}(t)\leq 1$. The set of such sequence is not empty, and a trivial example is $T(t)=t+\delta$ for some parameter $\delta>0$, which satisfies 1) and is clearly independent of any of the system components. Moreover, the exponent $b^{2}\,(C_{\mathrm{ref}}-2\,C_{p})\,t-2\,b^{2}\,C_{\mathrm{ref}}^{-}\,T(t)$ is always negative under such an example.
Note that the factor $N^{-c}$ is likely suboptimal since it can not exceed the importance sampling factor $r$. This is intuitive, because we are essentially asking $N$ to compensate the SDE accumulation error. It is possible to relax from this and choose $c\geq r$. This in turn means that we need more assumptions on $C_{e}$ to ensure that $\overline{C}_{e}(t,T(t))$ is non-increasing in $t$, in contrast to what we aimed here for being increasing which is mild. However, such settings will make Assumption 2 more difficult to verify in reality. Since making mild and practically checkable assumptions on score approximation remains an open discussion (see, e.g., De Bortoli, 2022; Chen et al., 2023a), we opt for a suboptimal bound under assumptions that are simple and easy-to-verify (n.b., we have only assumed $C_{e}$ to be a positive non-increasing function without specifying any rate).
The result essentially reveals how we can leverage information from $\pi_{\mathrm{ref}}$ to achieve for lower resampling error. Commonly used resampling schemes assume that we only have access to the given samples $\{(w_{i},X_{i})\}_{i=1}^{N}$. On the contrary, diffusion resampling additionally assumes a sampler for $\pi_{\mathrm{ref}}$ that approximates $\pi$ yielding more information of the target. The diffusion resampling effectively fuses the samples from $\{(w_{i},X_{i})\}_{i=1}^{N}$ and $\pi_{\mathrm{ref}}$ to obtain the re-samples, with the constant $b$ indicating how much we rely on $\pi_{\mathrm{ref}}$. With $b=0$, it means that we completely trust the samples from $\pi_{\mathrm{ref}}$, and this is indeed optimal when $\pi_{\mathrm{ref}}=\pi$. See Appendix D for more details.
@@ Appendix E Common experiment settings
Implementations are based on JAX (Bradbury et al., 2018). When applying a Wasserstein distance to quantify the quality of approximate samples, we use the earth-moving cost function denoted by $\mathsf{W}_{1}$, approximated by a sliced Wasserstein distance with 1,000 projections. We directly use the implementation by OTT-JAX (Cuturi et al., 2022). For neural network implementations, we use Flax (Heek et al., 2024).
@@
  • Four SDE integrators. The commonly used Euler–Maruyama, and the two exponential integrators by Jentzen and Kloeden (2009) and Lord and Rougemont (2004), see Section 2.3 for formulae, and Tweedie’s formula (Ho et al., 2020, essentially DDPM). They are called by EM, Jentzen–Kloeden, Lord–Rougemont, and Tweedie, respectively in the later context.
  • SDE and probability flow. The resampling SDE in Equation 4 can be solved also with a probability flow ODE (Song et al., 2021). We test both SDE and ODE versions.
@@
The seminal work by Jang et al. (2017) did not formulate how the Gumbel trick can be applied for resampling. We here make it explicit. The gist is to approximate the discrete indexing of multinomial resampling by a matrix-vector multiplication which is differentiable. Recall our weighted samples $\{(w_{i},X_{i})\}_{i=1}^{N}$, and define a matrix $\mathbf{X}\in\mathbb{R}^{N\times d}$ aggregating all the $N$ samples. For each $i$, draw $N$ independent samples $\{u_{i,j}\sim\mathrm{Uniform}[0,1]\}_{j=1}^{N}$ and compute $g_{i,j}=-\log\log u_{i,j}$. Then, the $i$-th re-sample is
@@
where $S_{i}\in\mathbb{R}^{1\times N}$ is a Softmax vector with elements $S_{i,j}=\overline{S}_{i,j}\,/\,\sum_{j=1}^{N}\overline{S}_{i,j}$, where $\overline{S}_{i,j}=\exp((\log w_{j}+g_{i,j})\,/\,\tau)$. This is independently repeated for $i=1,2,\ldots,N$ to generate the re-samples $\{X^{*}_{i}\}_{i=1}^{N}$. Upon the limit $\tau\to 0$, this recovers multinomial resampling, but the variance of the gradient estimation will grow too. See also Rosato et al. (2022).
@@
The approach by Karkus et al. (2018) view the samples $\{(w_{i},X_{i})\}_{i=1}^{N}$ as a discrete distribution with probability given by their normalised weights. Denote this discrete distribution by $\pi^{\mathrm{D}}$. Directly sampling from $\pi^{\mathrm{D}}$ using weighted random choice will result in undefined gradients. Soft resampling mitigates this issue by introducing another proposal (discrete) distribution $q^{\mathrm{D}}$ defined by $q^{\mathrm{D}}(X_{i})\coloneqq\alpha\,w_{i}+(1-\alpha)\,\frac{1}{N}$ for $i=1,2,\ldots,N$. Then, sampling from $\pi^{\mathrm{D}}$ can be achieved by an importance sampling as follows.